# Towards Improving Learning from Demonstration Algorithms via MCMC Methods

### v2

@@ 3.2 Imitation Learning
We consider the framework of Markov Decision Processes (MDP). We assume an infinite horizon setting. At any given state $s\in{\mathcal{S}}$, an agent makes decisions via a stochastic policy $\pi:{\mathcal{S}}\times{\mathcal{A}}\to\mathbb{R}_{\geq 0}$. We denote a trajectory to be a sequence of state-action pairs $\tau=(s_{0},a_{0},s_{1},a_{1},\cdots)$. Any policy $\pi$, along with MDP parameters, induces a distribution over trajectories, which can be expressed as $p_{\pi}(\tau)=p(s_{0})\prod_{t=0}^{\infty}\pi(a_{t}|s_{t}){\mathcal{T}}(s_{t+1}|s_{t},a_{t})$. The return of a trajectory is the discounted sum of rewards $R(\tau)=\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})$.
In imitation learning, we sidestep the availability of the reward function. Instead, we have access to a finite set of $D$ trajectories $\tau_{E}$ (a.k.a. demonstrations) that are sampled from an expert policy $\pi_{E}$. Every trajectory $\tau\in\tau_{E}$ consists of a finite-length sequence of state and action pairs $\tau=(s_{0},a_{0},s_{1},a_{1},\cdots)$, where $s_{0}\sim p_{\rm{0}}(s)$, $a_{t}\sim\pi_{E}(\cdot|s_{t})$, and $s_{t+1}\sim{\mathcal{T}}(\cdot|s_{t},a_{t})$. Our goal is to learn a parameterized policy $\pi_{\theta}$ which best approximates the expert policy given access to $\tau_{E}$.
@@ 3.3 Behavioral Cloning
Behavioral cloning (BC) casts imitation learning as a supervised learning problem over state-action pairs provided in the expert demonstrations. In particular, we learn the policy by solving a regression problem with states $s_{t}$ and actions $a_{t}$ as the features and target labels respectively. Formally, we minimize the following objective:
@@ 4.1 Generate Demonstrations by Trajectory Optimization
To generate demonstrations for rolling, we leverage a gradient-based trajectory optimization with the differentiable simulator. We write our trajectory optimization in Direct Shooting form Tedrake (2022) with all the constraints captured by the dynamics of the environment. Thus, given the initial state $s_{0}$, the objective is:
@@
where $s_{t},s_{g}$ are the current state and goal state, $d$ measures the similarity between the current state and the goal state, $L^{contact}$ encourages the robot to make contact to the object, and $\lambda$ weighs the different loss terms.
With differentiable dynamics, we can compute the gradient of the loss w.r.t the actions as follows:
@@
where both $L^{roll}_{t}$ and $L^{contact}_{t}$ are differentiable w.r.t. the ground-truth state $s_{t}$, and $\frac{\partial s_{t^{\prime}}}{\partial a_{t}}$ can be computed efficiently via the back-propagation technique using the differentiable dynamics model. Therefore, a straightforward approach used by many others Huang et al. (2021); Li et al. (2019) is to use gradient-descent to directly update the actions, as shown in Figure 1.

### v3

@@ 3.2 Imitation Learning
In imitation learning, we assume access to a set of demonstration trajectories trajectories $\tau_{E}\in D$ from an expert. Every trajectory $\tau=(s_{0},a_{0},s_{1},a_{1},\cdots)$ consists of a finite-length sequence of state and action pairs The goal of imitation learning is to learn a policy $\pi_{\theta}$ that imitates the expert policy used to generate the demonstration.
@@ 3.3 Behavioral Cloning
Behavioral cloning (BC) directly leverages supervised learning over state-action pairs in the expert demonstrations. In the deterministic case, the policy is learned by solving a regression problem with states $s_{t}$ as the features and actions $a_{t}$ and target labels. Formally, we minimize the following objective:
@@ 4.1 Generate Demonstrations by Trajectory Optimization
We leverage gradient-based trajectory optimization with a differentiable simulator to generate demonstration data. In Direct Shooting form Tedrake (2022), our trajectory optimization with all the constraints captured by the dynamics of the environment is shown below. Given the initial state $s_{0}$, the objective is:
@@
where $s_{t}$, is the current state, $s_{g}$ is the goal state, $d$ is the similarity between the two, $L^{contact}$ is a contact loss, and $\lambda$ weighs the different loss terms. With differentiable dynamics, we can compute the gradient of the loss with respect to the actions:
@@
Here, $\frac{\partial s_{t^{\prime}}}{\partial a_{t}}$ can be computed using the back-propagation technique with a differentiable dynamics model. Therefore, a straightforward approach used by many prior works Huang et al. (2021); Li et al. (2019); qi2022dough is to use gradient-descent to directly update the actions, as shown in Figure 1.